OpenStudy (anonymous):

what type of graph represents tan(x/2)=y?

OpenStudy (experimentx):

you know the graph of y = tan x .. right??

OpenStudy (anonymous):

yes is it any different? what would the change be?

OpenStudy (anonymous):

?

OpenStudy (experimentx):

the graph of this function ...or curve, is shrieked by half along x axis

OpenStudy (anonymous):

The period is changed. Normally the period is π. y = a tan (bx - c) + d y = a tan (x/2) The period is found by using π/b π/1/2 = 2π

OpenStudy (experimentx):

Let \[ f(x) = \tan x \\ \text{ what would you expect from } f(\frac x 2) ?? \]

OpenStudy (experimentx):

rest it just has same property as tan

OpenStudy (anonymous):

so basically could i do x over two times 2pi which would equal 4pi and however many passes on the graph are 4pi away is the answer?

OpenStudy (experimentx):

ah yes, if you do f(2x) ,, the period will double, if you do f(x/3) ,,, the period will reduce by 1/3

OpenStudy (anonymous):

so since its x/2 the graph will reduce but will the curve lines in the graph still curve the same way as the original graph? just be closer

OpenStudy (anonymous):

Look closely on how spaced out the units are. Normally, it's supposed to be measured in units of radians, but my software doesn't have that. \

OpenStudy (experimentx):

Oh ... sorry, my mistake ... i said the opposite

OpenStudy (anonymous):

omg thank you guys so much it helped now a stupid question lol how do i post another question because im new to this website and it doesnt show me where i can put in a question

OpenStudy (experimentx):

i forgot that x's have opposite relation with ... dividing the domain actually expands the curve along x-axis ... while multiplying shrinks

OpenStudy (anonymous):

TO post another question, close this question. THere should be a button by your question.

OpenStudy (anonymous):

ok thank you! :)

OpenStudy (anonymous):

You're welcome :)